OpenStudy (anonymous):

x/5 - x/2 = 22 How do I find the value of x?

hero (hero):

\(\large\frac{x}{5} - \frac{x}{2} = 22\)

OpenStudy (anonymous):

first find the lcd then cross multiply

OpenStudy (cwrw238):

2x - 5x = 220 -3x = 220